Chapter 17 .5 4. Calculate the pH a solution made with 25mL of a solution after being titrated with 0.25M HCI in the following amounts a. OmL 12 b. 10mL T c. 38mL 5,04 d. 50mL .40 I have the correct answers written on this questions, but I need help understanding how to get those answers. Thanks!
Answer #1

given moles of NH3-molarity Xvolume 0.38Mx0.025L-0.0095mol a) 10.38M E 0.38-х NH 0.38-x 0.38-x x is very small so that 0.38-x

b) after addition of 10 mL HCl moles of HCI added = 0.25M×0.01L=0.0025mol when 0.0025mol of HCl is added to the above base, s

C) after additi on of 38 mL HC moles of HCI = 0.2514× 0.038= 0.0095mol when 0.0095moles of HCl is added, 0.0095mol of NH, con